Find the equation of the line in slope intercept form that passes through (-3,-2) and is parallel to 2x-5y=20

Answer by Cromlix(4381)

First of all convert 2x - 5y =20
into y = mx + c form.
2x - 5y =20
-5y = - 2x + 20
y = 2/5 - 4
So new line being parallel to
the line above has the same gradient.
Secondly use y - b = m(x - a)
y + 2 = 2/5(x + 3)
5y + 10 = 2x + 6
5y = 2x + 6 -10
5y + 2x - 4